\begin{thebibliography}{10}

\bibitem{ANSTREICHER}
Kurt~M. Anstreicher.
\newblock The thirteen spheres: a new proof.
\newblock {\em Discrete Comput. Geom.}, 31(4):613--625, 2004.

\bibitem{BACHOCVELLENTIN}
Christine Bachoc and Frank Vallentin.
\newblock New upper bounds for kissing numbers from semidefinite programming.
\newblock {\em J. Amer. Math. Soc.}, 21(3):909--924, 2008.

\bibitem{BACHOCVALLENTIN2}
Christine Bachoc and Frank Vallentin.
\newblock Semidefinite programming, multivariate orthogonal polynomials, and
  codes in spherical caps.
\newblock {\em European J. Combin.}, 30(3):625--637, 2009.

\bibitem{BANNAIBANNAI}
Eiichi Bannai and Etsuko Bannai.
\newblock A survey on spherical designs and algebraic combinatorics on spheres.
\newblock {\em European J. Combin.}, 30(6):1392--1425, 2009.

\bibitem{BANNAISLOANE}
Eiichi Bannai and N.~J.~A. Sloane.
\newblock Uniqueness of certain spherical codes.
\newblock {\em Canadian J. Math.}, 33(2):437--449, 1981.

\bibitem{BARGMUSIN}
Alexander Barg and Oleg~R. Musin.
\newblock Codes in spherical caps.
\newblock {\em Adv. Math. Commun.}, 1(1):131--149, 2007.

\bibitem{BOROCZKY}
K\'{a}roly B\"{o}r\"{o}czky.
\newblock The {N}ewton-{G}regory problem revisited.
\newblock In {\em Discrete geometry}, volume 253, pages 103--110. Dekker, New
  York, 2003.

\bibitem{BOROCZKYGLAZYRIN}
K\'{a}roly~J. B\"{o}r\"{o}czky and Alexey Glazyrin.
\newblock Stability of optimal spherical codes.
\newblock {\em Monatsh. Math.}, 205(3):455--475, 2024.

\bibitem{BOYVALENKOVDRAGNEVHARDINSTOYANOVA2}
P.~G. Boyvalenkov, P.~D. Dragnev, D.~P. Hardin, E.~B. Saff, and M.~M.
  Stoyanova.
\newblock Universal lower bounds for potential energy of spherical codes.
\newblock {\em Constr. Approx.}, 44(3):385--415, 2016.

\bibitem{BOYVALENKOVDRAGNEVHARDINSTOYANOVA}
P.~G. Boyvalenkov, P.~D. Dragnev, D.~P. Hardin, E.~B. Saff, and M.~M.
  Stoyanova.
\newblock Bounds for spherical codes: the {L}evenshtein framework lifted.
\newblock {\em Math. Comp.}, 90(329):1323--1356, 2021.

\bibitem{BOYVALENKOVDANEV}
Peter Boyvalenkov and Danyo Danev.
\newblock On maximal codes in polynomial metric spaces.
\newblock In {\em Applied algebra, algebraic algorithms and error-correcting
  codes}, volume 1255 of {\em Lecture Notes in Comput. Sci.}, pages 29--38.
  Springer, Berlin, 1997.

\bibitem{BOYVALENKOVDANEVLANDGEV}
Peter Boyvalenkov, Danyo Danev, and Ivan Landgev.
\newblock On maximal spherical codes. {II}.
\newblock {\em J. Combin. Des.}, 7(5):316--326, 1999.

\bibitem{BOYVALENKOVDODUNEKOVMUSIN}
Peter Boyvalenkov, Stefan Dodunekov, and Oleg Musin.
\newblock A survey on the kissing numbers.
\newblock {\em Serdica Math. J.}, 38(4):507--522, 2012.

\bibitem{CASSELMAN}
Bill Casselman.
\newblock The difficulties of kissing in three dimensions.
\newblock {\em Notices Amer. Math. Soc.}, 51(8):884--885, 2004.

\bibitem{COHNJIAOKUMARTORQUATO}
Henry Cohn, Yang Jiao, Abhinav Kumar, and Salvatore Torquato.
\newblock Rigidity of spherical codes.
\newblock {\em Geom. Topol.}, 15(4):2235--2273, 2011.

\bibitem{COHNZHAO}
Henry Cohn and Yufei Zhao.
\newblock Sphere packing bounds via spherical codes.
\newblock {\em Duke Math. J.}, 163(10):1965--2002, 2014.

\bibitem{CONWAYSLOANE}
J.~H. Conway and N.~J.~A. Sloane.
\newblock {\em Sphere packings, lattices and groups}.
\newblock Springer-Verlag, New York, 1999.

\bibitem{DELSARTEGOETHALSSEIDEL}
P.~Delsarte, J.~M. Goethals, and J.~J. Seidel.
\newblock Spherical codes and designs.
\newblock {\em Geometriae Dedicata}, 6(3):363--388, 1977.

\bibitem{DIAGANABOOK}
Toka Diagana.
\newblock {\em Non-{A}rchimedean linear operators and applications}.
\newblock Nova Science Publishers, Inc., Huntington, NY, 2007.

\bibitem{DIAGANARAMAROSON}
Toka Diagana and Francois Ramaroson.
\newblock {\em Non-{A}rchimedean operator theory}.
\newblock SpringerBriefs in Mathematics. Springer, Cham, 2016.

\bibitem{ERICSONZINOVIEV}
Thomas Ericson and Victor Zinoviev.
\newblock {\em Codes on {E}uclidean spheres}, volume~63 of {\em North-Holland
  Mathematical Library}.
\newblock North-Holland Publishing Co., Amsterdam, 2001.

\bibitem{GLAZYRIN}
Alexey Glazyrin.
\newblock A short solution of the kissing number problem in dimension three.
\newblock {\em Discrete Comput. Geom.}, 69(3):931--935, 2023.

\bibitem{JENSSENJOOSPERKINS}
Matthew Jenssen, Felix Joos, and Will Perkins.
\newblock On kissing numbers and spherical codes in high dimensions.
\newblock {\em Adv. Math.}, 335:307--321, 2018.

\bibitem{KALISCH}
G.~K. Kalisch.
\newblock On {$p$}-adic {H}ilbert spaces.
\newblock {\em Ann. of Math. (2)}, 48:180--192, 1947.

\bibitem{KALLALKANWANG}
Kenz Kallal, Tomoka Kan, and Eric Wang.
\newblock Improved lower bounds for kissing numbers in dimensions 25 through
  31.
\newblock {\em SIAM J. Discrete Math.}, 31(3):1895--1908, 2017.

\bibitem{KUKLIN}
N.~A. Kuklin.
\newblock Delsarte method in the problem on kissing numbers in high-dimensional
  spaces.
\newblock {\em Proc. Steklov Inst. Math.}, 284(1):S108--S123, 2014.

\bibitem{LEECH}
John Leech.
\newblock The problem of the thirteen spheres.
\newblock {\em Math. Gaz.}, 40:22--23, 1956.

\bibitem{LIBERTI}
Leo Liberti.
\newblock Mathematical programming bounds for kissing numbers.
\newblock In {\em Optimization and decision science: methodologies and
  applications}, volume 217 of {\em Springer Proc. Math. Stat.}, pages
  213--222. Springer, Cham, 2017.

\bibitem{MACHADOFILHO}
Fabr\'{\i}cio~Caluza Machado and Fernando~M\'{a}rio de~Oliveira~Filho.
\newblock Improving the semidefinite programming bound for the kissing number
  by exploiting polynomial symmetry.
\newblock {\em Exp. Math.}, 27(3):362--369, 2018.

\bibitem{MAEHARA}
H.~Maehara.
\newblock The problem of thirteen spheres---a proof for undergraduates.
\newblock {\em European J. Combin.}, 28(6):1770--1778, 2007.

\bibitem{MITTELMANNVALLENTIN}
Hans~D. Mittelmann and Frank Vallentin.
\newblock High-accuracy semidefinite programming bounds for kissing numbers.
\newblock {\em Experiment. Math.}, 19(2):175--179, 2010.

\bibitem{MUSIN3}
O.~R. Musin.
\newblock Bounds for codes by semidefinite programming.
\newblock {\em Tr. Mat. Inst. Steklova}, 263:143--158, 2008.

\bibitem{MUSIN1}
Oleg~R. Musin.
\newblock The kissing problem in three dimensions.
\newblock {\em Discrete Comput. Geom.}, 35(3):375--384, 2006.

\bibitem{MUSIN2}
Oleg~R. Musin.
\newblock The kissing number in four dimensions.
\newblock {\em Ann. of Math. (2)}, 168(1):1--32, 2008.

\bibitem{ODLYZKOSLOANE}
A.~M. Odlyzko and N.~J.~A. Sloane.
\newblock New bounds on the number of unit spheres that can touch a unit sphere
  in {$n$} dimensions.
\newblock {\em J. Combin. Theory Ser. A}, 26(2):210--214, 1979.

\bibitem{PFENDER}
Florian Pfender.
\newblock Improved {D}elsarte bounds for spherical codes in small dimensions.
\newblock {\em J. Combin. Theory Ser. A}, 114(6):1133--1147, 2007.

\bibitem{PFENDERZIEGLER}
Florian Pfender and G\"{u}nter~M. Ziegler.
\newblock Kissing numbers, sphere packings, and some unexpected proofs.
\newblock {\em Notices Amer. Math. Soc.}, 51(8):873--883, 2004.

\bibitem{SAMORODNITSKY}
Alex Samorodnitsky.
\newblock On linear programming bounds for spherical codes and designs.
\newblock {\em Discrete Comput. Geom.}, 31(3):385--394, 2004.

\bibitem{SARDARIZARGAR}
Naser~Talebizadeh Sardari and Masoud Zargar.
\newblock New upper bounds for spherical codes and packings.
\newblock {\em Math. Ann.}, 389(4):3653--3703, 2024.

\bibitem{SCHUTTEVANDERWAERDEN}
K.~Sch\"{u}tte and B.~L. van~der Waerden.
\newblock Das {P}roblem der dreizehn {K}ugeln.
\newblock {\em Math. Ann.}, 125:325--334, 1953.

\bibitem{SLOANE}
N.~J.~A. Sloane.
\newblock Tables of sphere packings and spherical codes.
\newblock {\em IEEE Trans. Inform. Theory}, 27(3):327--338, 1981.

\bibitem{ZONG}
Chuanming Zong.
\newblock {\em Sphere packings}.
\newblock Universitext. Springer-Verlag, New York, 1999.

\end{thebibliography}

\documentclass{amsart}
\usepackage{setspace}
\usepackage{a4}
\usepackage{amsthm}
\usepackage{latexsym}
\usepackage{amsfonts}
\usepackage{graphicx}
\usepackage{textcomp}
\usepackage{cite}
\usepackage{enumerate}
\usepackage{amssymb}
\usepackage{hyperref}
\usepackage{amsmath}
\usepackage{tikz}
\usepackage[mathscr]{euscript}
\usepackage{mathtools}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{corollary}[theorem] {Corollary}
\newtheorem{definition}[theorem]{Definition}
\newtheorem{example}[theorem]{Example}

\newtheorem{problem}[theorem]{Problem}

\title{This is the title}
\usepackage{fancyhdr}

\begin{document}
\hrule\hrule\hrule\hrule\hrule
\vspace{0.3cm}	
\begin{center}
{\bf{p-adic Delsarte-Goethals-Seidel-Kabatianskii-Levenshtein-Pfender Bound}}\\
\vspace{0.3cm}
\hrule\hrule\hrule\hrule\hrule
\vspace{0.3cm}
\textbf{K. Mahesh Krishna}\\
School of Mathematics and Natural Sciences\\
Chanakya University Global Campus\\
NH-648, Haraluru Village\\
Devanahalli Taluk, 	Bengaluru  Rural District\\
Karnataka State, 562 110, India \\
Email: kmaheshak@gmail.com

Date: \today
\end{center}

\hrule\hrule
\vspace{0.5cm}
\textbf{Abstract}: We introduce the notion of p-adic spherical codes (in particular, p-adic kissing number problem). We show that the  one-line proof for a variant of the Delsarte-Goethals-Seidel-Kabatianskii-Levenshtein upper bound for spherical codes, obtained by Pfender \textit{[J. Combin. Theory Ser. A, 2007]}, extends to p-adic Hilbert spaces.

\textbf{Keywords}:  Spherical code, Kissing number, Linear programming, p-adic Hilbert space.

\textbf{Mathematics Subject Classification (2020)}: 52C17, 12J25, 46S10, 11D88.\\

\hrule

\hrule
\section{Introduction}
 Let $d\in \mathbb{N}$ and $\theta \in [0, 2\pi)$. A set $\{\tau_j\}_{j=1}^n$  of unit vectors  in $\mathbb{R}^d$	is said to be \textbf{$(d,n,\theta )$-spherical code} \cite{ZONG} in $\mathbb{R}^d$ if 
 	\begin{align}\label{SCI}
 	\langle \tau_j, \tau_k\rangle\leq \cos \theta , \quad \forall 1\leq j, k \leq n, j \neq k.
 \end{align}
Since 
\begin{align*}
	\langle \tau, \omega\rangle =\frac{2-\|\tau-\omega\|^2}{2}, \quad \forall \tau, \omega \in \mathbb{R}^d, 
\end{align*}
we can rewrite Inequality (\ref{SCI}) as 
\begin{align*}\label{SCN}
	\|\tau_j-\tau_k\|\geq  \sqrt{2(1-\cos \theta)}, \quad \forall 1\leq j, k \leq n, j \neq k.	
\end{align*}	
 Fundamental problem associated with spherical codes is the following.
 \begin{problem}\label{SCP}
 	Given $d$ and $\theta$, what is the maximum $n$ such that there exists a $(d,n,\theta )$-spherical code $\{\tau_j\}_{j=1}^n$  in $\mathbb{R}^d$?
 \end{problem}
 The case $\theta=\pi/3$ is known as the famous \textbf{(Newton-Gregory) kissing number problem}. With extensive efforts from many mathematicians, it is still not completely resolved in every dimension (but resolved in dimensions $d=1$ ($n=2$), $d=2$ ($n=6$), $d=3$ ($n=12$), $d=4$ ($n=24$), $d=8$ ($n=240$), $d=24$ ($n=196560$)) \cite{ANSTREICHER, PFENDER, CASSELMAN, MUSIN1, MUSIN2, ODLYZKOSLOANE,  SCHUTTEVANDERWAERDEN,  PFENDERZIEGLER, BACHOCVELLENTIN, MITTELMANNVALLENTIN, BOYVALENKOVDODUNEKOVMUSIN, BOROCZKY, MAEHARA, GLAZYRIN, LIBERTI, LEECH, KALLALKANWANG, JENSSENJOOSPERKINS, MACHADOFILHO, KUKLIN}. We refer  \cite{DELSARTEGOETHALSSEIDEL, BOYVALENKOVDRAGNEVHARDINSTOYANOVA, BARGMUSIN, BOYVALENKOVDRAGNEVHARDINSTOYANOVA2, ERICSONZINOVIEV, SARDARIZARGAR, MUSIN3, BANNAISLOANE, BACHOCVALLENTIN2, CONWAYSLOANE, COHNJIAOKUMARTORQUATO, SAMORODNITSKY, BOYVALENKOVDANEV, BOYVALENKOVDANEVLANDGEV, SLOANE, BANNAIBANNAI, BOROCZKYGLAZYRIN}
  for more on spherical codes.
 Problem \ref{SCP} has connection even with sphere packing  \cite{COHNZHAO}. Most used method for obtaining upper bounds on spherical codes is the Delsarte-Goethals-Seidel-Kabatianskii-Levenshtein
bound which we recall. Let $n \in \mathbb{N}$ be fixed. The Gegenbauer polynomials are defined inductively as 
\begin{align*}
	G_0^{(n)}(r)&\coloneqq1,\quad \forall r \in [-1,1],\\
	G_1^{(n)}(r)&\coloneqq r,\quad \forall r \in [-1,1],\\
	&\quad\vdots\\
	G_k^{(n)}(r)&\coloneqq\frac{(2k+n-4)rG_{k-1}^{(n)}(r)-(k-1)	G_{k-2}^{(n)}(r)}{k+n-3}, \quad \forall r \in [-1,1], \quad \forall k \geq 2.
\end{align*}
Then the family $\{G_k^{(n)}\}_{k=0}^\infty$ is orthogonal on the interval $[-1,1] $ with respect to the weight 
\begin{align*}
	\rho(r)\coloneqq (1-r^2)^\frac{n-3}{2}, \quad \forall r \in [-1,1].
\end{align*}
\begin{theorem} \cite{DELSARTEGOETHALSSEIDEL, ERICSONZINOVIEV} (\textbf{Delsarte-Goethals-Seidel-Kabatianskii-Levenshtein Linear Programming Bound}) \label{DGS}
	Let $\{\tau_j\}_{j=1}^n$  be a  $(d,n,\theta )$-spherical code in $\mathbb{R}^d$. 
	Let $P$ be a real polynomial satisfying following conditions.
	\begin{enumerate}[\upshape(i)]
		\item $P(r)\leq 0$ for all $-1\leq r\leq \cos \theta$.
		\item Coefficients in the  Gegenbauer expansion 
		\begin{align*}
			P=\sum_{k=0}^{m}a_kG_k^{(n)}
		\end{align*}
	satisfy 
	\begin{align*}
a_0>0, \quad a_k\geq 0, ~\forall 1\leq k \leq m.
	\end{align*}
	\end{enumerate}
Then
\begin{align*}
	n \leq \frac{P(1)}{a_0}.
\end{align*}
\end{theorem}
 In 2007, Pfender gave a one-line proof for a variant of Theorem  \ref{DGS}.
 \begin{theorem}  \cite{PFENDER} (\textbf{Delsarte-Goethals-Seidel-Kabatianskii-Levenshtein-Pfender Bound}) \label{PFENDERT}
 	Let $\{\tau_j\}_{j=1}^n$  be a  $(d,n,\theta )$-spherical code in $\mathbb{R}^d$. Let $c>0$ and $\phi:[-1,1]\to \mathbb{R}$ be a function satisfying following.
 	\begin{enumerate}[\upshape(i)]
 		\item 
 		\begin{align*}
 			\sum_{j=1}^n\sum_{k =1}^n\phi(\langle \tau_j, \tau_k\rangle)\geq 0.	
 		\end{align*}
 		\item $\phi(r)+c\leq 0$ for all $-1\leq r \leq \cos \theta$.
 	\end{enumerate}
 Then 
 \begin{align*}
 	n\leq \frac{\phi(1)+c}{c}.
 \end{align*}
In particular, if $\phi(1)+c\leq 1$, then $n\leq 1/c$.
 \end{theorem}
In this paper, we introduce the notion of p-adic spherical codes. We show that Theorem  \ref{PFENDERT},  can be easily extended for p-adic Hilbert spaces.

 \section{p-adic spherical Codes}
We begin from  the   definition  of  p-adic Hilbert space.  
\begin{definition}\cite{DIAGANABOOK, KALISCH, DIAGANARAMAROSON} \label{PADICDEF}
	Let $\mathbb{K}$ be a non-Archimedean complete valued field (with valuation $|\cdot|$) and $\mathcal{X}$ be a non-Archimedean Banach space (with norm $\|\cdot\|$) over $\mathbb{K}$. We say that $\mathcal{X}$ is a \textbf{p-adic Hilbert space} if there is a map (called as p-adic inner product) $\langle \cdot, \cdot \rangle: \mathcal{X} \times \mathcal{X} \to \mathbb{K}$ satisfying following.
	\begin{enumerate}[\upshape (i)]
		\item If $x \in \mathcal{X}$ is such that $\langle x,y \rangle =0$ for all $y \in \mathcal{X}$, then $x=0$.
		\item $\langle x, y \rangle =\langle y, x \rangle$ for all $x,y \in \mathcal{X}$.
		\item $\langle \alpha x+y, z \rangle =\alpha \langle x,  z \rangle+\langle y,z\rangle$ for all  $\alpha  \in \mathbb{K}$, for all $x,y,z \in \mathcal{X}$.
		\item $|\langle x, y \rangle |\leq \|x\|\|y\|$ for all $x,y \in \mathcal{X}$.
	\end{enumerate}
\end{definition}
Following is the  standard example which we consider in the paper.
\begin{example}\cite{KALISCH}
	Let $p$ be a prime. For $d \in \mathbb{N}$, let $\mathbb{Q}_p^d$ be the standard p-adic Hilbert space equipped with the inner product 
	\begin{align*}
		\langle (a_j)_{j=1}^d,(b_j)_{j=1}^d\rangle := \sum_{j=1}^da_jb_j,  \quad \forall (a_j)_{j=1}^d,(b_j)_{j=1}^d \in \mathbb{Q}_p^d
	\end{align*}
	and norm 
	\begin{align*}
		\|(x_j)_{j=1}^d\|:= \max_{1\leq j \leq d}|x_j|, \quad \forall (x_j)_{j=1}^d\in 	\mathbb{Q}_p^d.
	\end{align*}
\end{example}
We introduce p-adic spherical codes  as follows. 
\begin{definition}\label{PCODEDEFINITION}
	Let $d\in \mathbb{N}$ and $\theta \in [0, 2 \pi)$. A set $\{\tau_j\}_{j=1}^n$  of vectors  in $ \mathbb{Q}_p^d$	is said to be \textbf{p-adic $(d,n,\theta)$-spherical code}  in $\mathbb{Q}_p^d$ if following conditions hold.
	\begin{enumerate}[\upshape(i)]
		\item $	\|\tau_j\|=1$ for all  $1\leq j \leq n$.
		\item $	\langle \tau_j, \tau_j\rangle=1$ for all  $1\leq j \leq n$.
		\item  
		\begin{equation}\label{PE}
			|2-2\langle \tau_j, \tau_k\rangle|\geq 2(1-\cos \theta), \quad \forall 1\leq j, k \leq n, j \neq k.
		\end{equation}
		\end{enumerate}
		We call the  case $\theta=\pi/3$  as the \textbf{p-adic kissing number problem}.
\end{definition}
Let $\{\tau_j\}_{j=1}^n$  be a   p-adic $(d,n,\theta )$-spherical code  in $\mathbb{Q}_p^d$. 
Since 
\begin{align*}
	\|\tau_j-\tau_k\|^2\geq |\langle \tau_j-\tau_k,  \tau_j-\tau_k\rangle|=|2-2\langle \tau_j, \tau_k\rangle|, \quad  \forall 1\leq j, k \leq n, 
\end{align*}
Inequality   (\ref{PE}) gives
\begin{align}\label{PN}
	\|\tau_j-\tau_k\|\geq \sqrt{2(1-\cos \theta)},	\quad  \forall 1\leq j, k \leq n,  j \neq k.
\end{align}
However, note that Inequality (\ref{PN}) may not give Inequality   (\ref{PE}). Note that we can formulate the definition of  general p-adic $(d,n,\theta)$-spherical code by replacing Inequality (\ref{PE})  with Inequality (\ref{PN}) in Definition \ref{PCODEDEFINITION}. It is also possible to consider Definition \ref{PCODEDEFINITION} by removing conditions (i) or (ii).
Following is the p-adic version of Theorem \ref{PFENDERT}.
\begin{theorem} (\textbf{p-adic Delsarte-Goethals-Seidel-Kabatianskii-Levenshtein-Pfender Spherical Codes Bound}) \label{PDGSP}
	Let $\{\tau_j\}_{j=1}^n$  be a  p-adic $(d,n,\theta )$-spherical code in $\mathbb{Q}_p^d$. Let $c>0$ and $\phi:[0, \infty)\to \mathbb{R}$ be a function satisfying following.
	\begin{enumerate}[\upshape(i)]
		\item $\sum_{1\leq j,k \leq n}\phi(|2-2\langle \tau_j, \tau_k\rangle|)\geq 0$.
		\item $\phi(r)+c\leq 0$ for all $ r \in [2(1-\cos \theta), \infty)$.
	\end{enumerate}
	Then 
	\begin{align*}
		n\leq \frac{\phi(0)+c}{c}.
	\end{align*}
	In particular, if $\phi(0)+c\leq 1$, then $n\leq 1/c$.	
\end{theorem}
\begin{proof}
	Define $\psi:[0, \infty)\ni r \mapsto \psi(r)\coloneqq \phi(r)+c\in \mathbb{R}$. Then 
	\begin{align*}
		\sum_{1\leq j,k \leq n}\psi(|2-2\langle \tau_j, \tau_k\rangle|)&=\sum_{j=1}^{n}\psi(|2-2\langle \tau_j, \tau_j\rangle|)+\sum_{1\leq j,k \leq n, j \neq k}\psi(|2-2\langle \tau_j, \tau_k\rangle|)\\
		&=\sum_{j=1}^{n}\psi(0)+\sum_{1\leq j,k \leq n, j \neq k}\psi(|2-2\langle \tau_j, \tau_k\rangle|)\\
		&=n(\phi(0)+c)+\sum_{1\leq j,k \leq n, j \neq k}(\phi(|2-2\langle \tau_j, \tau_k\rangle|+c)\\
		&\leq n(\phi(0)+c)+0=n(\phi(0)+c).
	\end{align*}
	We also have 
	\begin{align*}
		\sum_{1\leq j,k \leq n}\psi(\phi(|2-2\langle \tau_j, \tau_k\rangle|))=\sum_{1\leq j,k \leq n}(\phi(|2-2\langle \tau_j, \tau_k\rangle|)+c)=\sum_{1\leq j,k \leq n}\phi(|2-2\langle \tau_j, \tau_k\rangle|)+cn^2.
	\end{align*}
	Therefore 
	\begin{align*}
		cn^2\leq \sum_{1\leq j,k \leq n}\phi(|2-2\langle \tau_j, \tau_k\rangle|)+cn^2\leq n(\phi(0)+c).
	\end{align*}
\end{proof}
\begin{corollary} (\textbf{p-adic Delsarte-Goethals-Seidel-Kabatianskii-Levenshtein-Pfender Kissing Number Bound})
	Let $\{\tau_j\}_{j=1}^n$  be a  p-adic $(d,n,\pi/3)$-spherical code in $\mathbb{Q}_p^d$. Let $c>0$ and $\phi:[0, \infty)\to \mathbb{R}$ be a function satisfying following.
	\begin{enumerate}[\upshape(i)]
		\item $\sum_{1\leq j,k \leq n}\phi(|2-2\langle \tau_j, \tau_k\rangle|)\geq 0$.
		\item $\phi(r)+c\leq 0$ for all $ r \in [1, \infty)$.
	\end{enumerate}
	Then 
	\begin{align*}
		n\leq \frac{\phi(0)+c}{c}.
	\end{align*}
	In particular, if $\phi(0)+c\leq 1$, then $n\leq 1/c$.	
\end{corollary}	
Following generalization of Theorem \ref{PDGSP} is easy.
\begin{theorem}
	Let $\{\tau_j\}_{j=1}^n$  be a  p-adic $(d,n,\theta )$-spherical code in $\mathbb{Q}_p^d$. Let $c>0$ and 
	\begin{align*}
		\phi:\{|2-2\langle \tau_j, \tau_k\rangle|:1\leq j, k \leq n\}\to \mathbb{R}	
	\end{align*}
	be a function satisfying following.
	\begin{enumerate}[\upshape(i)]
		\item $\sum_{1\leq j,k \leq n}\phi(|2-2\langle \tau_j, \tau_k\rangle|)\geq 0$.
		\item $\phi(r)+c\leq 0$ for all $ r \in \{|2-2\langle \tau_j, \tau_k\rangle|:1\leq j, k \leq n, j \neq k\}$.
	\end{enumerate}
	Then 
	\begin{align*}
		n\leq \frac{\phi(0)+c}{c}.
	\end{align*}
	In particular, if $\phi(0)+c\leq 1$, then $n\leq 1/c$.	
\end{theorem}
Note that, in the paper, we can replace $\mathbb{Q}_p^d$ by any p-adic Hilbert space.

 \bibliographystyle{plain}
 \bibliography{reference.bib}

\end{document}